Resource Usage and Supply, Activity Rates, Service Organization EcoBrite Labs performs tests on water samples supplied by outside companies to ensure that their waste water meets environmental standards. Customers deliver water samples to the lab and receive the lab reports via the Internet. The EcoBrite Labs facility is built and staffed to handle the processing of 100,000 tests per year. The lab facility cost $200,000 to build and is expected to last 10 years and will have no salvage value. Processing equipment cost $240,000 and has a life expectancy of five years and will have no salvage value. Both facility and equipment are depreciated on a straight-line basis. EcoBrite Labs has five salaried laboratory technicians, each of whom is paid $35,000. In addition to the salaries, facility, and equipment, EcoBrite Labs expects to spend $54,000 for chemicals and other supplies (assuming 100,000 tests are performed). Last year, 89,000 tests were performed.

Required:

2. Calculate the total annual activity rate for the water testing activity. Break the activity rate into fixed and variable components. (Round your answers to three significant digits.)

Total water testing rate: $ ____ per test

Variable activity rate: $_____ per test

Fixed activity rate: $_____ per test

3. Compute the total activity availability, and break this into activity output and unused activity.

Activity availability:_____ tests

Activity usage: ______tests

Unused activity:______ tests

4. Calculate the total cost of resources supplied, and break this into the cost of activity used and the cost of unused activity.

Cost of activity supplied: $ ____

Cost of activity used: $ ____

Cost of unused activity: $ _____
Answer #1

Part 2

Total water testing rate = facility cost+ Processing equipment cost+salary of laboratory technicians+supplies expense = ((200000/10)+(240000/5)+(35000*5)+54000)/100000 = $2.97 per test

Variable activity rate = supplies expense = 54000/100000 = $0.54 per test

Fixed activity rate = $2.97- $0.54 = $2.43 per test

Part 3

Activity availability = 100000 tests

Activity usage = 89000 tests

Unused activity = 11000 tests

Part 4

Cost of activity supplied = (200000/10)+(240000/5)+(35000*5)+(0.54*89000) = $291060

Cost of activity used = (2.97*89000) = 264330

Cost of unused activity = (2.43*11000) = $26730